# Model-Based Transfer Learning for Contextual Reinforcement Learning

### v2

@@ Abstract
Deep reinforcement learning (RL) is a powerful approach to complex decision making. However, one issue that limits its practical application is its brittleness, sometimes failing to train in the presence of small changes in the environment. Motivated by the success of zero-shot transfer—where pre-trained models perform well on related tasks—we consider the problem of selecting a good set of training tasks to maximize generalization performance across a range of tasks. Given the high cost of training, it is critical to select training tasks strategically, but not well understood how to do so. We hence introduce Model-Based Transfer Learning (MBTL), which layers on top of existing RL methods to effectively solve contextual RL problems. MBTL models the generalization performance in two parts: 1) the performance set point, modeled using Gaussian processes, and 2) performance loss (generalization gap), modeled as a linear function of contextual similarity. MBTL combines these two pieces of information within a Bayesian optimization (BO) framework to strategically select training tasks. We show theoretically that the method exhibits sublinear regret in the number of training tasks and discuss conditions to further tighten regret bounds. We experimentally validate our methods using urban traffic and standard continuous control benchmarks. The experimental results suggest that MBTL can achieve up to 50x improved sample efficiency compared with canonical independent training and multi-task training. Further experiments demonstrate the efficacy of BO and the insensitivity to the underlying RL algorithm and hyperparameters. This work lays the foundations for investigating explicit modeling of generalization, thereby enabling principled yet effective methods for contextual RL. Code is available at https://github.com/jhoon-cho/MBTL/.
@@ 1 Introduction
We build upon zero-shot transfer, a widely-used practical technique which directly applies a policy trained in one context (source task) to another (target task) without adaptation. This experiences performance degradation, also called generalization gap [17, 22]. For instance, Figure 1 depicts that the performance degrades as the target task diverges from the source task, corresponding to an increasing generalization gap. Nonetheless, leveraging the notion that training is expensive but zero-shot transfer is cheap, we are interested in optimally selecting a set of source (training) tasks, such that the generalization performance on the target range of tasks is maximized. In this article, we strategically select source tasks by explicitly modeling the generalization performance to estimate the value of training a new source task.
@@
  • We introduce Model-Based Transfer Learning (MBTL), a novel framework for studying generalization in reinforcement learning (Figure 2). To the best of our knowledge, this is the first work to explicitly model generalization performance for contextual RL. As such, our work opens the door for further investigation into reliable model-based methodologies for contextual RL.
@@
  • We empirically validate our methods in urban traffic and standard continuous control benchmarks for contextual RL, observing up to 50x improvements in sample efficiency. We further include ablations on the components of the algorithm.
@@ 2 Preliminaries and notation
Generalization gap via zero-shot transfer. Consider zero-shot transfer from the trained policy $\pi_{x}$ from a source task (context-MDP) to solve another target task (context-MDP) with the context $x^{\prime}\in X$ in the CMDP. Zero-shot transfer involves applying a policy trained on a source task ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}_{x}}$ to a different target task ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}_{x^{\prime}}}$, with $x,x^{\prime}\in X$. We observe the generalization performance, denoted by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J(\pi_{x},x^{\prime}})$, by evaluating the target task $x^{\prime}$ based on the policy trained using source task $x$ via zero-shot generalization. We define the generalization gap as the absolute performance difference in average reward when transferring from source task $x$ to target task $x^{\prime}$:
$\underbrace{\Delta J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})}_{\text{Generalization gap}}=\underbrace{J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x)}_{\text{Source task performance}}-\underbrace{J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})}_{\text{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Generalization} performance}}.$ (1)
@@ 4 Model-Based Transfer Learning (MBTL)
In this section, we introduce an algorithm called Model-based Transfer Learning to solve the SSTS problem. MBTL models the generalization performance in two parts: 1) the performance set point, modeled using Gaussian processes, and 2) performance loss (generalization gap), modeled as a linear function of contextual similarity. MBTL combines these two pieces of information within a Bayesian optimization (BO) framework to sequentially select training tasks that maximize generalization performance. The high-level pseudo-code of the MBTL algorithm is given below.
MBTL Pseudo-code 1:Input: Task set, Gaussian Process (GP) model, and training budget 2:while training budget remains do 3:    Select next source task $\leftarrow$ Task with the highest acquisition value based on estimated training performance, generalization gap, and generalization performance, determined by the current GP model and tailored acquisition function 4:    Train a policy on the selected source task using reinforcement learning, then evaluate its training performance 5:    Update the GP model with the selected source task training performance to estimate the training performance of tasks in the task set 6:end while 7:Output: Set of policies that maximizes generalized performance across task set
@@ 4.1 Modeling assumptions
We consider a task set $X$ that is continuous and the performance function $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi},x),V(x)$ for a policy $\pi$ to be smooth over the task space $X$. In practice, such as control systems, tasks often vary continuously and smoothly rather than abruptly. For example, adjusting the angle of a robotic arm by a small amount typically results in a small change in the system and optimal action.
Inspired by the empirical generalization gap performance as observed in Figure 1, we estimate the generalization gap $\Delta\hat{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x^{\prime})$ with a linear function.
@@ 4.3 Regret analysis
We use regret to quantify the effectiveness of our source task selection based on BO. Specifically, we define regret at iteration $k$ as $r_{k}=\hat{V}(x_{k}^{*};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})-\hat{V}(x_{k};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$, where $\hat{V}(x_{k}^{*};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$ represents the maximum generalization performance achievable across all tasks, and $\hat{V}(x_{k};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$ is the performance at the current task selection $x_{k}$. Consequently, the cumulative regret after $K$ iterations is given by $R_{K}=\sum_{k=1}^{K}r_{k}$, summing the individual regrets over all iterations. Following the framework presented by Srinivas et al. [40], our goal is to establish that this cumulative regret grows sublinearly with respect to the number of iterations. Mathematically, we aim to prove that $\lim_{K\rightarrow\infty}\frac{R_{K}}{K}=0$, indicating that, on average, the performance of our strategy approaches the optimal performance as the number of iterations increases.
@@ 5.1 Setup
DRL algorithms and performance measure. We utilize Deep Q-Networks (DQN) for discrete action spaces [29] and Proximal Policy Optimization (PPO) for continuous action spaces [35]. We evaluate our method by the average performance across all $N$ target tasks after training up to $K=$15 source tasks or the number of source tasks needed to achieve a certain level of performance. We employ min-max normalization of the rewards for each task, and we provide comprehensive details about our model in Appendix A.4.1.
@@ 5.2 Traffic benchmark experiments
Table 1: Comparative performance of different methods on traffic CMDPs
Benchmark (CMDP) Baselines Sequential Baselines MBTL Oracle
@@
Traffic Signal Inflow 0.8646 0.8319 0.8457 0.8700 0.8496 0.8673 0.8768
Traffic Signal Speed Limit 0.8857 0.6083 0.8821 0.8858 0.8862 0.8854 0.8876
Eco-driving Penetration Rate 0.5260 0.1945 0.5959 0.5934 0.5827 0.6323 0.6660
Eco-driving Inflow 0.4061 0.2229 0.4774 0.4705 0.4673 0.5108 0.5528
Eco-driving Green Phase 0.3850 0.4228 0.4406 0.4557 0.4431 0.4700 0.5027
AA-Ring-Acc Hold Duration 0.8362 0.9209 0.8924 0.9057 0.8776 0.9242 0.9552
AA-Ring-Vel Hold Duration 0.9589 0.972 0.9785 0.9772 0.9807 0.9816 0.9822
AA-Ramp-Acc Hold Duration 0.4276 0.5158 0.6050 0.5956 0.6143 0.6318 0.7111
AA-Ramp-Vel Hold Duration 0.5473 0.5034 0.669 0.6787 0.5907 0.7182 0.7686
Average 0.6778 0.6017 0.7312 0.7354 0.722 0.7559 0.7844
@@
We consider three traffic control domains. First, while most traffic lights still operate on fixed schedules, RL can be used to design adaptive (1) Traffic signal control to optimize traffic [9, 23]. However, considering that every intersection is different, challenges persist in generalizing across intersection configurations [19]. Given the significant portion of greenhouse gas emissions in the United States due to transportation [12], the second traffic domain is (2) Dynamic eco-driving at signalized intersections, which concerns learning energy-efficient driving strategies in urban settings. DRL-based eco-driving strategies have been developed [14, 45, 18] but still experience difficulties in generalization. Our final traffic domain is (3) Advisory autonomy, in which real-time speed or acceleration advisories guide human drivers to act as vehicle-based traffic controllers in mixed traffic environments [39, 8, 16]. The context space $X$ is discretized into $N=\{50,50,40\}$ contexts for the three domains, respectively. In Appendix A.4, we provide details about our experiments.
Figure 4: Traffic CMDP results. Method comparison of normalized performance over $N$ tasks. MBTL efficiently selects source training tasks. The black dotted line indicates the first training step within MBTL that exceeds both independent and multi-task baselines, with up to 25x fewer samples needed.
Results. Table 1 and Figure 4 summarize the results. Notably, the Oracle far outperforms the standard baselines (Independent and Multi-task), indicating the potential for transfer learning and intelligent training of multiple models, respectively. MBTL rapidly approaches the Oracle, within $\approx 10$ transfer steps, indicating that the Gaussian process effectively models the training performance and linear generalization gap models generalization performance. It is important to note that multi-task RL studies commonly consider Independent training as a strong baseline due to its avoidance of negative transfer and other training instability issues. Indeed, Independent often (but not always) outperforms Multi-task in our experiments. Yet, our experiments show that MBTL consistently outperforms both Independent and Multi-task and matches their performance with 5-25x better sample efficiency. Among the sequential baselines, MBTL consistently outperforms the heuristic baselines, indicating the value of adaptively selecting source tasks based on feedback. The sequential baselines, such as random, equidistant, and greedy strategy, also generally outperform Independent and Multi-task, indicating the general value of multi-policy training for solving CMDPs. More results are provided in Appendix A.4.

### v3

@@ Abstract
Deep reinforcement learning (RL) is a powerful approach to complex decision making. However, one issue that limits its practical application is its brittleness, sometimes failing to train in the presence of small changes in the environment. Motivated by the success of zero-shot transfer—where pre-trained models perform well on related tasks—we consider the problem of selecting a good set of training tasks to maximize generalization performance across a range of tasks. Given the high cost of training, it is critical to select training tasks strategically, but not well understood how to do so. We hence introduce Model-Based Transfer Learning (MBTL), which layers on top of existing RL methods to effectively solve contextual RL problems. MBTL models the generalization performance in two parts: 1) the performance set point, modeled using Gaussian processes, and 2) performance loss (generalization gap), modeled as a linear function of contextual similarity. MBTL combines these two pieces of information within a Bayesian optimization (BO) framework to strategically select training tasks. We show theoretically that the method exhibits sublinear regret in the number of training tasks and discuss conditions to further tighten regret bounds. We experimentally validate our methods using urban traffic and standard continuous control benchmarks. The experimental results suggest that MBTL can achieve up to 43x improved sample efficiency compared with canonical independent training and multi-task training. Further experiments demonstrate the efficacy of BO and the insensitivity to the underlying RL algorithm and hyperparameters. This work lays the foundations for investigating explicit modeling of generalization, thereby enabling principled yet effective methods for contextual RL. Code is available at https://github.com/jhoon-cho/MBTL/.
@@ 1 Introduction
Figure 1: Normalized performance comparison across different problem variations in Eco-Driving benchmark. Traditional DRL approaches (e.g., Independent training or multi-task training) exhibit greater training instability, whereas Oracle Transfer, zero-shot transfer with full information, shows the potential for performance improvement by multi-policy training.
We build upon zero-shot transfer, a widely-used practical technique which directly applies a policy trained in one context (source task) to another (target task) without adaptation. Figure 1 shows that multi-policy training with zero-shot transfer has the potential to improve the performance even over the independent training. In this article, we strategically select source tasks by explicitly modeling the generalization performance to estimate the value of training a new source task.
@@
  • We introduce Model-Based Transfer Learning (MBTL), a novel framework for solving CMDP sample efficiently (Figure 2). To the best of our knowledge, this is the first work to explicitly model generalization performance for contextual RL (CRL). As such, our work opens the door for further investigation into reliable model-based methodologies for CRL.
@@
  • We empirically validate our methods in urban traffic and standard continuous control benchmarks for contextual RL, observing up to 43x improvements in sample efficiency. We further include ablations on the components of the algorithm.
@@ 2 Preliminaries and notation
Generalization gap via zero-shot transfer. Consider zero-shot transfer from the trained policy $\pi_{x}$ from a source task (context-MDP) to solve another target task (context-MDP) with the context $x^{\prime}\in X$ in the CMDP. Zero-shot transfer involves applying a policy trained on a source task ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}_{x}}$ to a different target task ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}_{x^{\prime}}}$, with $x,x^{\prime}\in X$. This experiences performance degradation, also called generalization gap [17, 23]. For instance, Figure 3 depicts that the performance degrades as the target task diverges from the source task, corresponding to an increasing generalization gap. Nonetheless, leveraging the notion that training is expensive but zero-shot transfer is cheap, we are interested in optimally selecting a set of source (training) tasks, such that the generalization performance on the target range of tasks is maximized. We observe the generalization performance, denoted by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J(\pi_{x},x^{\prime}})$, by evaluating the target task $x^{\prime}$ based on the policy trained using source task $x$ via zero-shot generalization. We define the generalization gap as the absolute performance difference in average reward when transferring from source task $x$ to target task $x^{\prime}$:
$\underbrace{\Delta J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})}_{\text{Generalization gap}}=|\underbrace{J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x)}_{\text{Source task performance}}-\underbrace{J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})}_{\text{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Generalization} performance}}|.$ (1)
@@ 4 Model-Based Transfer Learning (MBTL)
In this section, we introduce an algorithm called Model-based Transfer Learning to solve the SSTS problem. MBTL models the generalization performance in two parts: 1) the performance set point, modeled using Gaussian processes, and 2) generalization gap, modeled as a linear function of contextual similarity. MBTL combines these two pieces of information within a Bayesian optimization (BO) framework to sequentially select training tasks that maximize generalization performance.
@@ 4.1 Modeling assumptions
We consider a task set $X$ that is continuous and the performance function $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi},x),V(x)$ for a policy $\pi$ to be smooth over the task space $X$. In practice, such as control systems, tasks often vary continuously and smoothly rather than abruptly. For example, adjusting the angle of a robotic arm by a small amount typically results in a small change in the system and optimal action. Inspired by the empirical generalization gap performance as observed in Figure 3, we estimate the generalization gap with a linear function of contextual similarity.
@@ 4.3 Regret analysis
We use regret to quantify the effectiveness of our source task selection based on BO. Specifically, we define regret at iteration $k$ as $r_{k}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}(x_{k}^{*};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}(x_{k};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$, where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V}(x_{k}^{*};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$ represents the maximum generalization performance achievable across all tasks, and $\hat{V}(x_{k};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})$ is the performance at the current task selection $x_{k}$. Consequently, the cumulative regret after $K$ iterations is given by $R_{K}=\sum_{k=1}^{K}r_{k}$, summing the individual regrets over all iterations. Following the framework presented by Srinivas et al. [41], our goal is to establish that this cumulative regret grows sublinearly with respect to the number of iterations. Mathematically, we aim to prove that $\lim_{K\rightarrow\infty}\frac{R_{K}}{K}=0$, indicating that, on average, the performance of our strategy approaches the optimal performance as the number of iterations increases.
@@ 5.1 Setup
DRL algorithms and performance measure. We utilize Deep Q-Networks (DQN) for discrete action spaces [29] and Proximal Policy Optimization (PPO) for continuous action spaces [36]. For statistical reliability, we run each experiment three times with different random seeds. We evaluate our method by the average performance across all $N$ target tasks after training up to $K=$15 source tasks or the number of source tasks needed to achieve a certain level of performance. We employ min-max normalization of the rewards for each task, and we provide comprehensive details about our model in Appendix A.4.1.
@@ 5.2 Traffic benchmark experiments
Table 1: Comparative performance of different methods on traffic CMDPs ($K=15$)
Benchmark (CMDP) Baselines Multi-policy Baselines MBTL Oracle
@@
Traffic Signal Road Length 0.9409 0.8242 0.9366 0.9337 0.9349 0.9364 0.9409
Traffic Signal Inflow 0.8646 0.8319 0.8699 0.8712 0.8682 0.8432 0.8768
Traffic Signal Speed Limit 0.8857 0.6083 0.8872 0.8872 0.8874 0.8867 0.8876
Eco-Driving Penetration Rate 0.526 0.1945 0.6212 0.6278 0.5992 0.6148 0.666
Eco-Driving Inflow 0.4061 0.2229 0.5077 0.5513 0.5299 0.5172 0.5528
Eco-Driving Green Phase 0.385 0.4228 0.4724 0.4697 0.4678 0.4985 0.5027
AA-Ring-Acc Hold Duration 0.8362 0.9209 0.9306 0.9225 0.9136 0.9382 0.9552
AA-Ring-Vel Hold Duration 0.9589 0.972 0.9819 0.9819 0.9819 0.9818 0.9822
AA-Ramp-Acc Hold Duration 0.4276 0.5158 0.6750 0.6821 0.6882 0.6964 0.7111
AA-Ramp-Vel Hold Duration 0.5473 0.5034 0.7170 0.7386 0.6918 0.7273 0.7686
Average 0.6778 0.6017 0.7600 0.7666 0.7563 0.7641 0.7844
@@
We consider three traffic control benchmarks. First, while most traffic lights still operate on fixed schedules, RL can be used to design adaptive (1) Traffic signal control to optimize traffic [8, 24]. However, considering that every intersection is different, challenges persist in generalizing across intersection configurations [19]. Given the significant portion of greenhouse gas emissions in the United States due to transportation [11], the second traffic domain is (2) Dynamic eco-driving at signalized intersections, which concerns learning energy-efficient driving strategies in urban settings. DRL-based eco-driving strategies have been developed [13, 47, 18] but still experience difficulties in generalization. Our final traffic domain is (3) Advisory autonomy, in which real-time speed or acceleration advisories guide human drivers to act as vehicle-based traffic controllers in mixed traffic environments [40, 7, 15]. The context space $X$ is discretized into $N=\{50,50,40\}$ contexts for the three domains, respectively. In Appendix A.4, we provide details about our experiments.
Figure 5: Traffic CMDP results. Method comparison of normalized performance over $N$ tasks. MBTL efficiently selects source training tasks. The black dotted line indicates the first training step within MBTL that exceeds both independent and multi-task baselines, with up to 30x fewer samples.
Results. Table 1 and Figure 5 summarize the results. Notably, the Oracle far outperforms the standard baselines (independent and multi-task training), indicating the potential for transfer learning and intelligent training of multiple models, respectively. MBTL rapidly approaches the Oracle within $\approx 10$ transfer steps, indicating that the GP effectively models the training performance and linear generalization gap models the generalization performance. It is important to note that multi-task RL studies commonly consider Independent training as a strong baseline due to its avoidance of negative transfer and other training instability issues. Indeed, independent training often (but not always) outperforms multi-task training in our experiments. Yet, our experiments show that MBTL outperforms both independent and multi-task baselines and matches their performance with up to 30x improved sample efficiency. Among the multi-policy baselines, MBTL often outperforms the heuristic multi-policy baselines, indicating the value of adaptively selecting source tasks based on feedback. The multi-policy baselines, such as random, equidistant, and greedy strategy, also generally outperform Independent and Multi-task, indicating the general value of multi-policy training for solving CMDPs. More results are provided in Appendix A.4.